OpenStudy (anonymous):

Determine the degree of the product. -2x^2(4x^3-5x^2)

OpenStudy (anonymous):

2+3 = 5

OpenStudy (anonymous):

-2x^2(4x^3-5x^2) = -8x^5 +10x^4

OpenStudy (anonymous):

degree of the product. is nothing but highest deress of any trem -2x^2(4x^3-5x^2) = -8x^5 +10x^4 so the highest degree is 5